Based on the concept on testing the difference between two means using z-test and the learning exercises that you have done, write your arguments or lessons learned below.


1.

2.

3.

4.

5.


1
Expert's answer
2021-10-20T14:25:06-0400

Z-test is a statistical test that compares the difference between two population means. Two random samples of sizes "n_1" and "n_2" respectively are drawn from two normal populations with means "\mu_1" and "\mu_2" and variances "\sigma^2_1" and "\sigma^2_2".

Conditions for this test to be performed include,

  • Two samples must be randomly selected.
  • The samples must be independent.
  • Each sample size must be at least 30 and the population variances are known.

If these requirements are met, the difference of the two sample means follow a standard normal distribution with mean "\mu_{\bar{x}_1-\bar{x}_2}=\mu_1-\mu_2" and variance, "\sigma^2_{\bar{x}_1-\bar{x}_2}=\sqrt{(\sigma_1^2/n_1)+(\sigma^2_2/n_2)}".

The Null hypothesis to be tested is given as, "H_0:\mu_1=\mu_2" against the alternative hypothesis that can either be two-sided or one-sided and they can be written as below.

Two sided alternative hypothesis

"H_1:\mu_1\not=\mu_2"

One sided alternative hypothesis

"H_1:\mu_1\gt\mu_2" or "\mu_1\lt\mu_2"

The test statistic used is,

"Z_0=(\bar{x}_1-\bar{x}_2)/\sqrt{(\sigma_1^2/n_1)+(\sigma^2_2/n_2)}" '

The null hypothesis is rejected if the calculated "Z_0" is such that "|Z_0|\gt Z_{\alpha/2}" for a two sided test and for a one sided test it is rejected if "Z_0\gt Z_\alpha" for upper tailed one sided test and if "Z_0\lt-Z_\alpha" for lower tailed one sided test for a specified "\alpha" level of significance.